(a) According to the Ministry of Health, the height of Guinean travellers who were quarantined in Tamale for the novel coronavirus were normally distributed about a mean of 160cm and a standard deviation of 8cm. Find the probability that a traveller selected at random has
i) Height between 148cm and 175cm
ii) Height above 164cm
iii) Height below 179cm
1
Expert's answer
2020-06-03T19:53:52-0400

We denote by a random variable, which is a height of a randomly selected traveller. It has a normal distribution with "\mu=160" cm, "\sigma=8" cm. We remind that its probability density function has the form "p(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}" . We obtain the following probabilities (rounded to 4 decimal places):

i) "P(148\leq X\leq175)=\frac{1}{8\sqrt{2\pi}}\int_{148}^{175}e^{-\frac12(\frac{x-160}{8})^2}dx\approx0.9028" ;

ii) "P(X\geq164)=\frac{1}{8\sqrt{2\pi}}\int_{164}^{+\infty}e^{-\frac12(\frac{x-160}{8})^2}dx\approx0.3085";

iii) "P(X\leq179)=\frac{1}{8\sqrt{2\pi}}\int_{-\infty}^{179}e^{-\frac12(\frac{x-160}{8})^2}dx\approx0.9912".


Anaconda (https://www.anaconda.com/) was used for computations. The following code produces numeric values:


from scipy import integrate

import numpy as np

import math


func = lambda x:(1/(8*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-160)/8)*((x-160)/8))

Pr1 = integrate.quad(func, 148, 175)

Pr2 = integrate.quad(func, 164, np.Infinity)

Pr3 = integrate.quad(func, -np.Infinity, 179)


print(Pr1)

print(Pr2)

print(Pr3)


Answer: 0.9028; 0.3085; 0.9912 (rounded to 4 decimal places)